# Efficient Generation of Multi-partite Entanglement between Non-local Superconducting Qubits using Classical Feedback

### v2

@@ Abstract
Quantum entanglement is one of the primary features which distinguishes quantum computers from classical computers. In gate-based quantum computing, the creation of entangled states or the distribution of entanglement across a quantum processor often requires circuit depths which grow with the number of entangled qubits. However, in teleportation-based quantum computing, one can deterministically generate entangled states with a circuit depth that is constant in the number of qubits, provided that one has access to an entangled resource state, the ability to perform mid-circuit measurements, and can rapidly transmit classical information. In this work, aided by fast classical FPGA-based control hardware with a feedback latency of only 150 ns, we explore the utility of teleportation-based protocols for generating non-local, multi-partite entanglement between superconducting qubits. First, we demonstrate well-known protocols for generating Greenberger-Horne-Zeilinger (GHZ) states and non-local CNOT gates in constant depth. Next, we utilize both protocols for implementing a quantum fan-out gate in constant depth among three non-local qubits (i.e., controlled-NOT-NOT). Finally, we demonstrate deterministic state teleportation and entanglement swapping between qubits on opposite sides of our quantum processor.
@@ I Introduction
Quantum entanglement plays a central role in nearly all quantum applications. In gate-based quantum computing, entanglement is generated by multi-qubit unitary operations (i.e., gates), which can be used to prepare multi-partite entangled states. In systems with all-to-all connectivity, generating multi-partite entanglement can be trivially accomplished by means of two-qubit gates Moses et al. (2023). However, in systems with limited connectivity (e.g., superconducting circuits), generating non-local multi-partite entanglement by unitary operators can be costly due to the need for SWAP networks for swapping states between neighboring qubits O’Gorman et al. (2019); Hashim et al. (2022). However, long-range entangled states can be efficiently prepared using quantum teleportation-based protocols Bennett et al. (1993); Gottesman and Chuang (1999) in a hardware-agnostic manner. Quantum teleportation is a method for transmitting a quantum state from one qubit to another, provided that they each share half of a Bell state and can transmit classical information. When embedded within a larger (unitary) quantum circuit for performing adaptive control of entangled states, teleportation protocols necessitate the use of mid-circuit measurements (MCMs) and fast classical feed-forward operations conditioned on the results of the measured outcomes. Using adaptive circuits, it has been shown that no purely unitary circuit can prepare the same entangled state for the same circuit depth and connectivity Foss-Feig et al. (2023). Moreover, these adaptive circuits have been proposed for simulating exotic phases of matter with non-Abelian anyons Bravyi et al. (2022); Tantivasadakarn et al. (2022a, b), long-range interacting quantum matter Lu et al. (2022, 2023), and for preparing topologically ordered states Iqbal et al. (2023). Finally, as the size of quantum computers continues to grow, adopting modular computing architectures Monroe et al. (2014) will be necessary. In such architectures, quantum teleportation protocols Gottesman and Chuang (1999); Bennett et al. (1993); Jiang et al. (2007) will be useful tools for transmitting quantum information from one system to another.
@@
In the noisy intermediate-scale quantum (NISQ) era Preskill (2018), qubit coherence times and gate fidelities limit the circuit depth with which we can perform useful computations. For superconducting qubits, if long-range entanglement is needed between qubits which are not physically connected, swapping information requires circuit depth for qubits Hashim et al. (2022). Therefore, generating multi-partite entanglement between non-local qubits becomes intractable if the available size of contemporary systems grows faster than the fidelity of multi-qubit gates can be improved. However, it has been shown that non-local, multi-partite entangled states can be prepared in constant depth Briegel and Raussendorf (2001); Foss-Feig et al. (2023); Moses et al. (2023); Bäumer et al. (2023) using teleportation-based protocols, as long as one has access to mid-circuit measurements and adaptive feedback. These entangled states can further be used as a resource for other teleportation-based protocols (e.g., generating non-local entangling gates).
@@
In this work, we utilize open source FPGA-based control hardware with low feedback latency (150 ns) Xu et al. (2021, 2023a); Fruitwala et al. (2024) for exploring the utility of teleportation-based protocols on an eight-qubit superconducting quantum processor (see Appendix). We first demonstrate well-known procedures for preparing Greenberger–Horne–Zeilinger (GHZ) states and implementing non-local teleportation-based CNOT gates in constant depth. Then, we show how one can utilize both protocols to implement an unbounded fan-out gate in constant depth, which we use to create a controlled-NOT-NOT gate between three unconnected qubits. Finally, we show how one can deterministically teleport quantum states or generate Bell pairs between distant qubits in constant-depth, and demonstrate these protocols between qubits on opposite sides of our quantum processor. Throughout this work, we explore the extent to which MCMs dephase idling spectator qubits, and find that the measurement-induced dephasing Gambetta et al. (2006) strongly depends on the spectral resolution of the readout resonators coupled to each qubit, which limits the size and complexity of the protocols that we can implement on our current processor. This work serves as a useful study for understanding how teleportation-based protocols can be utilized as a resource within larger (unitary-based) quantum circuits, and what limitations need to be overcome for utilizing such protocols for superconducting qubits at scale.
@@ II GHZ State Preparation
The first task we demonstrate is the preparation of a GHZ state between non-nearest neighbor qubits in constant depth. GHZ states are highly entangled states that play an important role in quantum foundations Greenberger et al. (1989); Mermin (1990); Caves et al. (2002), quantum metrology Eldredge et al. (2018), and fault-tolerant quantum computation Shor (1996). To prepare an -qubit GHZ state with one- and two-qubit gates requires at least circuit depth even with all-to-all connectivity Watts et al. (2019). However, if we are allowed to perform MCMs and implement gate operations on the unmeasured qubits based on the classical outcomes, the depth can be reduced to constant regardless of the size of the GHZ state. The idea for constant depth GHZ state preparation can be easily understood via the stabilizer formalism. The GHZ state is the eigenstate for a set of stabilizer generators with eigenvalue . Therefore, one can initialize all data qubits in state and use ancillae qubits in the middle of two data qubits to measure type stabilizers in parallel. If all the measurement outcomes are , then the state among those data qubits should be in the standard form of the GHZ state, provided that there is no readout error. On the other hand, if some of the measurement outcomes are , this indicates that there exist some “domain walls” between adjacent data qubits and therefore we need an extra layer of single-qubit gates acting on certain qubits to recover the standard GHZ state. While this protocol has already been demonstrated in trapped ions Moses et al. (2023), cold atoms Bluvstein et al. (2024), and superconducting qubits Bäumer et al. (2023), we similarly show it here as it not only serves as the backbone for other tasks like the constant-depth fan-out gate implementation, but also provides one figure of merit to benchmark the performance of measurement- and teleportation-based protocols in our platform.
@@
The circuit schematic for preparing a four-qubit GHZ state in constant depth on our eight-qubit processor is shown in Fig. 1a. Here, four data qubits (black) are entangled with three ancillae qubits (red) in an alternating fashion. By performing an MCM of the ancillae qubits, we can learn about the parity of each pair of data qubits, and the results of the parity measurements can be decoded in real-time to determine which data qubit(s) experienced a bit-flip (see Appendix D). In Figs. 1b–d, we plot the raw (marginalized) outputs of measurements made in the computational basis for two-, three-, and four-qubit GHZ states prepared among data qubits. While the classical outputs resemble GHZ states (albeit very noisy in the four-qubit case), we use parity oscillations Sackett et al. (2000); Leibfried et al. (2005); Monz et al. (2011) to verify whether or not we have generated genuine entanglement among our data qubits. The fidelity of an -qubit GHZ state state can subsequently be computed via
@@
where is the population of the state, is the population of the state, and is the amplitude of oscillations of the measured parity (i.e., the coherence of the state). If , this represents a case of genuine entanglement between -qubits Leibfried et al. (2005); Dür and Cirac (2001); Gühne and Seevinck (2010). For our constant-depth creation of GHZ states, we measure a fidelity of , , and for the two-, three-, and four-qubit GHZ states, respectively. Therefore, while we observe genuine entanglement for the two- and three-qubit GHZ states, the four-qubit GHZ fails the above criteria.
@@
In general, the fidelity of states prepared via teleportation- and measurement-based protocols can be limited by many sources of errors, including readout errors ( decay, crosstalk, etc.), noise on idling spectator qubits, gate errors, etc. In our case, the errors are likely due to measurement-induced dephasing on spectator qubits (see Appendix E). Because superconducting qubits are measured via dispersive coupling to a readout resonator, depending on the frequency separation of the resonators of neighboring qubits and the static coupling between nearest neighbors, the measurement of an ancilla qubit can result in the dephasing of a spectator qubit. If the dephasing is weak, it can be mitigated via dynamical decoupling. However, if the readout resonator of the ancilla qubit is close in frequency to the readout resonator of the data qubit, then measurement of the ancilla qubit can indirectly measure the data qubit, in which case phase coherence of the data qubit is completely lost after the MCM (we discuss this in detail in Appendix E). This dephasing can be mitigated by ensuring that ancilla qubits are always those whose resonators are spectrally well-resolved from the resonators of the data qubits, but this can sometimes necessitate SWAP gates between qubits, which increases the contribution from gate errors in the preparation circuit.
@@ III Teleportation-based CNOT
Teleportation-based quantum gates are an important component of quantum internet and networking protocols Kimble (2008); Muralidharan et al. (2016); Jiang et al. (2009). Moreover, such gates enable entanglement between non-local qubits on systems with sparse connectivity, which can ease topology constraints when frequency crowding is an issue Morvan et al. (2022). It is also beneficial in the realm of quantum error correction. In contrast to the standard surface code Fowler et al. (2012) where only one logical qubit is encoded in each patch, quantum low-density parity check (QLDPC) codes Breuckmann and Eberhardt (2021) benefit from higher encoding rates, but come at the expense of non-local stabilizer checks. Therefore, it becomes difficult to perform stabilizer measurements for QLDPC codes when the connectivity of the qubit layout is limited by locality. Several hardware-efficient protocols have been proposed to tackle this problem. For example, neutral atoms trapped in optical tweezers can be moved around to perform entangling gates with other atoms far from each other Xu et al. (2023b); Bluvstein et al. (2024). On the other hand, for systems without all-to-all coupling (e.g., superconducting circuits), a teleportation-based CNOT is another option if entanglement is shared between two separate parties, where the long-range GHZ states may serve as the desired entanglement resource. Teleportation-based CNOT gates have been previously demonstrated in optical systems Huang et al. (2004), trapped ions Wan et al. (2019), superconducting cavities Chou et al. (2018), and superconducting qubits Bäumer et al. (2023). In our work, we demonstrate a teleportation-based CNOT gate as it serves as a precursor to the more expressive gate, the unbounded fan-out (see next section), which can also be demonstrated in constant depth using similar methods.
@@
The circuit diagram for the teleportation-based CNOT is shown in Fig. 2a, where it can be seen that a CNOT between distant qubits (blue) can be implemented by means of two MCMs with feed-forward, as long as the two ancillae qubits (black) are prepared in a Bell state. If the ancillae qubits themselves are non-local, then we may first prepare them in a Bell state using the same procedure introduced in the previous section for the constant-depth GHZ state preparation (Fig. 2b). As a result, the ability to perform constant depth long-range GHZ preparation provides the opportunity to execute constant depth teleportation-based CNOT, regardless of the distance between two data qubits. Because we are limited by measurement-induced dephasing on spectator qubits (see Appendix E), we cannot characterize teleportation-based CNOT gates with high fidelity using standard methods, such as quantum process tomography. Instead, we perform truth table tomography, which is a partial tomography technique that is limited to input states prepared in the computational basis Hashim et al. (2024). In Figs. 2c–d, we plot the truth table for performing a CNOT between qubits on opposite sides of our eight-qubit quantum processor, which is arranged in a ring geometry (and thus has only linear connectivity between nearest neighbors). We calculate a truth table fidelity for each pair of gates,
@@
where is the Hilbert space dimension for qubits, is the experimental truth table, and is the ideal truth table. For a CNOT between qubits 1 and 4 (which are only separated by two ancillae, and thus only require unitary preparation of a Bell state), we find a truth table fidelity of . For a CNOT between qubits 0 and 4 (which are separated by three ancillae, and thus can utilize the constant depth GHZ state prep shown in Fig. 2b), we find a truth table fidelity of .

### v3

@@ Abstract
Quantum entanglement is one of the primary features which distinguishes quantum computers from classical computers. In gate-based quantum computing, the creation of entangled states or the distribution of entanglement across a quantum processor often requires circuit depths which grow with the number of entangled qubits. However, in teleportation-based quantum computing, one can deterministically generate entangled states with a circuit depth that is constant in the number of qubits, provided that one has access to an entangled resource state, the ability to perform mid-circuit measurements, and can rapidly transmit classical information. In this work, aided by fast classical field programmable gate array-based control hardware with a feedback latency of only 150 ns, we explore the utility of teleportation-based protocols for generating non-local, multi-partite entanglement between superconducting qubits. First, we demonstrate well-known protocols for generating Greenberger-Horne-Zeilinger (GHZ) states and non-local CNOT gates in constant depth. Next, we utilize both protocols for implementing a quantum fan-out gate in constant depth among three non-local qubits (i.e., controlled-NOT-NOT). Finally, we demonstrate deterministic state teleportation and entanglement swapping between qubits on opposite sides of our quantum processor. Throughout this work, we find that the fidelity of our teleportation-based protocols is limited by measurement-induced dephasing on idling spectator qubits. Therefore, our work serves as a useful study of the current benefits and limitations of teleportation-based protocols on contemporary superconducting quantum processors.
@@ I Introduction
Quantum entanglement plays a central role in nearly all quantum applications. In gate-based quantum computing, entanglement is generated by multi-qubit unitary operations (i.e., gates), which can be used to prepare multi-partite entangled states. In systems with all-to-all connectivity, generating multi-partite entanglement can be trivially accomplished by means of two-qubit gates moses2023race. However, in systems with limited connectivity (e.g., superconducting circuits), generating non-local multi-partite entanglement by unitary operators can be costly due to the need for SWAP networks for swapping states between neighboring qubits o2019generalized; hashim2022optimized. However, long-range entangled states can be efficiently prepared using quantum teleportation-based protocols bennett1993teleporting; gottesman1999demonstrating in a hardware-agnostic manner. Quantum teleportation is a method for transmitting a quantum state from one qubit to another, provided that they each share half of a Bell state and can transmit classical information. When embedded within a larger (unitary) quantum circuit for performing adaptive control of entangled states, teleportation protocols necessitate the use of mid-circuit measurements (MCMs) and fast classical feed-forward operations conditioned on the results of the measured outcomes. Using adaptive circuits, it has been shown that no purely unitary circuit can prepare the same entangled state for the same circuit depth and connectivity foss2023experimental. For example, a variety of novel entangled states can be deterministically prepared with adaptive circuits in constant depth, including certain classes of matrix product statessmith_constant-depth_2024, non-Abelian or other exotic topologically ordered statesbravyi2022adaptive; tantivasadakarn2022hierarchy; tantivasadakarn2022shortest; lu2022measurement; lu2023mixed; verresen_efficiently_2022, and W or low-weight Dicke statesbuhrman_state_2024-1; niu2024ac as well. Moreover, adaptive circuits also provide a significant sampling overhead reduction in certifying multi-partite quantum statesgupta_few_2025. Finally, as the size of quantum computers continues to grow, adopting modular computing architectures monroe2014large will be necessary. In such architectures, quantum teleportation protocols gottesman1999demonstrating; bennett1993teleporting; jiang2007distributed will be useful tools for transmitting quantum information from one system to another.
@@
In the noisy intermediate-scale quantum (NISQ) era preskill2018quantum, qubit coherence times and gate fidelities limit the circuit depth with which we can perform useful computations. For superconducting qubits, if long-range entanglement is needed between qubits which are not physically connected, swapping information requires circuit depth for qubits hashim2022optimized. Therefore, generating multi-partite entanglement between non-local qubits becomes intractable if the available size of contemporary systems grows faster than the fidelity of multi-qubit gates can be improved. However, as mentioned above, non-local, multi-partite entangled states can be prepared in constant depth using teleportation-based protocols, as long as one has access to mid-circuit measurements and adaptive feedback. These entangled states can further be used as a resource for other teleportation-based protocols (e.g., generating non-local entangling gates).
@@
In this work, we utilize open-source control hardware based on field programmable gate arrays (FPGAs) with low feedback latency (150 ns) xu2021qubic; xu2023qubic; fruitwala2024distributed for exploring the utility of teleportation-based protocols on an eight-qubit superconducting quantum processor (see Appendix). We first demonstrate well-known procedures for preparing Greenberger–Horne–Zeilinger (GHZ) states and implementing non-local teleportation-based controlled-NOT (CNOT) gates in constant depth. Then, we show how one can utilize both protocols to implement an unbounded fan-out gate in constant depth, which we use to create a controlled-NOT-NOT (CNOT-NOT) gate between three unconnected qubits. Finally, we show how one can deterministically teleport quantum states or generate Bell pairs between distant qubits in constant-depth, and demonstrate these protocols between qubits on opposite sides of our quantum processor. Throughout this work, we explore the extent to which MCMs dephase idling spectator qubits, and find that the measurement-induced dephasing gambetta2006qubit strongly depends on the spectral resolution of the readout resonators coupled to each qubit, which limits the size and complexity of the protocols that we can implement on our current processor. This work serves as a useful study for understanding how teleportation-based protocols can be utilized as a resource within larger (unitary-based) quantum circuits, and what limitations need to be overcome for utilizing such protocols for superconducting qubits at scale. We believe that adaptive circuits will become a useful tool for state-preparation and non-local gates when the circuit depths of the equivalent unitary circuits are prohibitive on systems with limited connectivity.
@@ II GHZ State Preparation
The first task we demonstrate is the preparation of a GHZ state between non-nearest neighbor qubits in constant depth. GHZ states are highly entangled states that play an important role in quantum foundations greenberger1989going; mermin1990quantum; caves2002unknown, quantum metrology eldredge2018optimal, and fault-tolerant quantum computation Shor1996. To prepare an -qubit GHZ state with one- and two-qubit gates requires at least circuit depth even with all-to-all connectivity watts_exponential_2019. However, if we are allowed to perform MCMs and implement gate operations on the unmeasured qubits based on the classical outcomes, the depth can be reduced to constant regardless of the size of the GHZ state. The idea for constant depth GHZ state preparation can be easily understood via the stabilizer formalism. The GHZ state is the eigenstate for a set of stabilizer generators with eigenvalue . Therefore, one can initialize all data qubits in state and use ancillae qubits in the middle of two data qubits to measure type stabilizers in parallel. If all the measurement outcomes are , then the state among those data qubits should be in the standard form of the GHZ state, provided that there is no readout error. On the other hand, if some of the measurement outcomes are , this indicates that there exist some domain walls 111The term domain wall here indicates the location between two sequences of qubits with different parities (qubits in each of the two sequences have the same parity) between adjacent data qubits and therefore we need an extra layer of single-qubit gates acting on certain qubits to recover the standard GHZ state. While this protocol has already been demonstrated in trapped ions moses2023race, cold atoms bluvstein2024logical, and superconducting qubits baumer2023efficient, we similarly show it here as it not only serves as the backbone for other tasks like the constant-depth fan-out gate implementation, but also provides one figure of merit to benchmark the performance of measurement- and teleportation-based protocols in our platform.
@@
The circuit schematic for preparing a four-qubit GHZ state in constant depth on our eight-qubit processor is shown in Fig. 1(a). Here, four data qubits (black) are entangled with three ancillae qubits (red) in an alternating fashion. By performing an MCM of the ancillae qubits, we can learn about the parity of each pair of data qubits, and the results of the parity measurements can be decoded in real-time to determine which data qubit(s) experienced a bit-flip (see Appendix D). In Figs. 1(b)–(d), we plot the raw (marginalized) outputs of measurements made in the computational basis for two-, three-, and four-qubit GHZ states prepared among data qubits. While the classical outputs resemble GHZ states (albeit very noisy in the four-qubit case), we use parity oscillations sackett2000experimental; leibfried2005creation; monz201114 to verify whether or not we have generated genuine entanglement among our data qubits. The fidelity of an -qubit GHZ state state can subsequently be computed via
@@
where is the population of the state, is the population of the state, and is the amplitude of oscillations of the measured parity (i.e., the coherence of the state) 222Equation 1 is quantitatively similar to the state fidelity of a GHZ state, such as what would be estimated using quantum state tomography hashim2024practical.. If , this represents a case of genuine entanglement between qubits leibfried2005creation; dur2001multiparticle; guhne2010separability. For our constant-depth creation of GHZ states, we measure a fidelity of , , and for the two-, three-, and four-qubit GHZ states, respectively. Therefore, while we observe genuine entanglement for the two- and three-qubit GHZ states, the four-qubit GHZ fails the above criteria.
@@
In general, the fidelity of states prepared via teleportation- and measurement-based protocols can be limited by many sources of errors. In Figs. 1(h)–(j), we plot error budgets for the two-, three-, and four-qubit GHZ state preparation experiments constructed using a back-of-the-envelope calculation from the various error rates and decoherence times listed in Appendix A. We broadly divide the error budgets into gate errors from single-qubit gates (“1Q Gates”), gate errors from two-qubit gates (“2Q Gates”), errors due to mid-circuit measurements and feed-forward operations (“MCM + FF”), decoherence due to decay, and dephasing333We perform readout correction on the classical output distributions, so we do not include errors due to terminating measurements in our error budgets. For example, we observe that as we increase the number of qubits in the experiment, dephasing errors grow (both in absolute value and relative to the rest of the error sources) to dominate the error budget. In our case, the dephasing errors are likely due to measurement-induced dephasing on spectator qubits (see Appendix E). Because superconducting qubits are measured via dispersive coupling to a readout resonator, depending on the frequency separation of the resonators of neighboring qubits and the static coupling between nearest neighbors, the measurement of an ancilla qubit can result in the dephasing of a spectator qubit. If the dephasing is weak, it can be mitigated via dynamical decoupling. However, if the readout resonator of the ancilla qubit is close in frequency to the readout resonator of the data qubit, then measurement of the ancilla qubit can indirectly measure the data qubit, in which case phase coherence of the data qubit is completely lost after the MCM (we discuss this in detail in Appendix E). This dephasing can be mitigated by ensuring that ancilla qubits are always those whose resonators are spectrally well resolved from the resonators of the data qubits, but this can sometimes necessitate SWAP gates between qubits, which increases the contribution from gate errors in the preparation circuit.
As noted above, the task of preparing GHZ states in constant depth using adaptive circuits has already been demonstrated in a number of different hardware platforms moses2023race; bluvstein2024logical; baumer2023efficient. In superconducting qubits, Ref. baumer2023efficient observes genuine entanglement in up to 8 qubits using adaptive circuits, genuine entanglement in up 13 qubits using just MCMs and updating the Pauli frame in post-processing, and genuine entanglement in up to 17 qubits using unitary circuits; all three versions fall short of state-of-the-art fidelities for GHZ states in superconducting qubits mooney2021generation. They highlight that their adaptive-based protocols were mainly limited by idling errors during MCM and feed-forward, compared to gate errors in unitary circuits. Thus, while they were able to successfully deploy the protocol on more qubits than we have demonstrated above, they did not show any improvement over unitary-based GHZ state preparation, limiting the utility of the protocol on contemporary superconducting quantum processors. In trapped ions, Ref. moses2023race successfully prepares a 32-qubit GHZ state using adaptive circuits with a fidelity of 0.74(1), but found that this did not outperform the unitary-based GHZ state [which had a fidelity of 0.82(1)] due to the fact that the adaptive circuit actually required more two-qubit gates and measurements than the equivalent unitary circuit; thus, in systems with all-to-all connectivity, adaptive circuits may not provide any benefit over unitary circuits. Finally, in cold atoms, Ref. bluvstein2024logical reports creating a three-qubit logical GHZ state using adaptive circuits with a fidelity of 0.77(2). They note that this fidelity is mainly limited by the underlying GHZ state preparations of the physical qubits, since atomic-based systems require shuttling their atoms between entangling zones, storage zones, and readout zones.
@@ III Teleportation-based CNOT
Teleportation-based quantum gates are an important component of quantum internet and networking protocols kimble2008quantum; muralidharan2016optimal; jiang2009quantum. Moreover, such gates enable entanglement between non-local qubits on systems with sparse connectivity, which can ease topology constraints when frequency crowding is an issue morvan2022optimizing. They may also be beneficial in the realm of quantum error correction. For example, for superconducting quantum processors, qubit connectivity is generally limited by locality. Therefore, for certain quantum error correcting codes that require nonlocal stabilizer measurements, the teleportation-based CNOT provides a low-depth solution for non-local entanglementdelfosse_bounds_2021. Teleportation-based CNOT gates have been previously demonstrated in optical systems huang2004experimental, trapped ions wan2019quantum, superconducting cavities chou2018deterministic, and superconducting qubits baumer2023efficient. In our work, we demonstrate a teleportation-based CNOT gate as it serves as a precursor to the more expressive gate, the unbounded fan-out (see next section), which can also be demonstrated in constant depth using similar methods.
@@
The circuit diagram for the teleportation-based CNOT is shown in Fig. 2(a), where it can be seen that a CNOT between distant qubits (blue) can be implemented by means of two MCMs with feed-forward, as long as the two ancillae qubits (black) are prepared in a Bell state. If the ancillae qubits themselves are non-local, then we may first prepare them in a Bell state using the same procedure introduced in the previous section for the constant-depth GHZ state preparation [Fig. 2(b)]. As a result, the ability to perform constant depth long-range GHZ preparation provides the opportunity to execute constant depth teleportation-based CNOT, regardless of the distance between two data qubits. Because we are limited by measurement-induced dephasing on spectator qubits (see Appendix E), we cannot characterize teleportation-based CNOT gates with high fidelity using standard methods, such as quantum process tomography. Instead, we perform truth table tomography, which is a partial tomography technique that is limited to input states prepared in the computational basis hashim2024practical. In Figs. 2(c)–(d), we plot the truth table for performing a CNOT between qubits on opposite sides of our eight-qubit quantum processor, which is arranged in a ring geometry (and thus has only linear connectivity between nearest neighbors). We calculate a truth table fidelity for each pair of gates,
@@
where is the Hilbert space dimension for qubits, is the experimental truth table, and is the ideal truth table 444Equation 2 is quantitatively similar to the process fidelity of a CNOT gate, such as what would be estimated using quantum process tomography hashim2024practical. In fact, a measurement of the truth table in two different bases can be used to lower bound the process fidelity hofmann2005complementary. For a CNOT between qubits 1 and 4 (which are only separated by two ancillae, and thus only require unitary preparation of a Bell state), we find a truth table fidelity of . For a CNOT between qubits 0 and 4 [which are separated by three ancillae, and thus can utilize the constant depth GHZ state prep shown in Fig. 2(b)], we find a truth table fidelity of .
In Figs. 2(e) and (f), we plot the error budgets for the CNOTs between qubits 1 and 4, and 0 and 4, respectively. For the CNOT between qubits 1 and 4, we do not observe any dephasing errors — this is due to the fact that the qubits remain in the computational basis for truth table tomography; therefore, they are insensitive to any measurement-induced dephasing. In contrast, for the CNOT between qubits 0 and 4, we observe some dephasing due to the fact that qubits 1 and 3 idle during the MCM used for the initial Bell state prep sub-circuit. In general, these fidelities and error budgets are only representative of the performance of the gate for state preparations and measurements in the computational basis; if we were able to prepare and measure in orthogonal bases, the effects of the measurement-induced dephasing would likely result in worse performance (see Appendix E) . We leave an exploration of this for future work.
Similar to the GHZ state preparation protocol, the teleportation-based CNOT is a well-known protocol bennett1993teleporting that has previously been implemented on various different hardware platforms huang2004experimental; chou2018deterministic; wan2019quantum; baumer2023efficient. Relevant to this work, Ref. baumer2023efficient finds that the teleportation-based CNOT outperforms the equivalent unitary circuit for 10 superconducting qubits in a linear chain. This improvement largely comes from reduction in the number of CNOT gates needed for the teleportation circuit compared to the unitary circuit, which requires decomposing SWAP gates into CNOTs. Moreover, they find that the performance of real-time correction is approximately equal to the equivalent measurement-based protocol with postprocessing up to about 40 qubits. This demonstrates that teleportation-based CNOTs can help alleviate topology constraints on devices with limited connectivity even on contemporary NISQ processors, effectively enabling all-to-all connectivity (as long as ancilla qubits are available).